OpenStudy (anonymous):

The width of a rectangle is 12 units. Can the perimeter x of the rectangle can be 60 units when its length y is 18 units? A:No, the rectangle cannot have x = 60 and y = 18 because x + y is less than 60 B:No, the rectangle cannot have x = 60 and y = 18 because x + y is less than 24 C:Yes, the rectangle can have x = 60 and y = 18 because x = 12 + 2y D:Yes, the rectangle can have x = 60 and y = 18 because x = 24 + 2y

OpenStudy (anonymous):

@perl @dan815 @wio

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@hughfuve

OpenStudy (anonymous):

@Hero @mathmath333 @perl @dan815 @wio

OpenStudy (anonymous):

@Mehek14 can you help me please

Mehek (mehek14):



Mehek (mehek14):

so add up 18+18+12+12 to find the perimeter

OpenStudy (anonymous):

60

Mehek (mehek14):

correct so the perimeter can be 60

Mehek (mehek14):

that means that you can eliminate A and B

OpenStudy (anonymous):

ok

Mehek (mehek14):

now the formula for perimeter is P = 2w + 2l w = width l = length

Mehek (mehek14):

in your case length is y so P = 2w + 2y

Mehek (mehek14):

you know that the width is 12 so P = 2(12) + 2y 2*12 = 24

Mehek (mehek14):

so your formula would be x = 24 + 2y

OpenStudy (anonymous):

So D

Mehek (mehek14):

yes

OpenStudy (anonymous):

Can you help with one more

Mehek (mehek14):

sure

OpenStudy (anonymous):

ill open a new question so u can get another medal

Mehek (mehek14):

alright thanks